Assuming that Coca Cola has the same specific heat as water (4.18 J/goC), calculate the amount of heat in kilojoules (kJ) transferred when one can (about 350g) is cooled from 25oC to 3oC.




1
Expert's answer
2020-11-02T09:05:50-0500

find the formula:

"Q=mc(T2-T1)=350\times4.18(3-25)=350\times4.18(-22)=-32186J=-32.186kJ"